# TextQuests: How Good are LLMs at Text-Based Video Games?

### v1

@@ 2 TextQuests
Previous work in text-based game evaluation has often relied on the games’ built-in scoring systems as the primary metric (Hausknecht et al., 2020; Yao et al., 2020). However, these point systems are a weak proxy for actual advancement, as they were often designed to reward exploration or enhance replayability rather than to track progress on the main storyline (for example, in The Witness, as many as 30 different endings are possible). To address these limitations, we introduce a new Game Progress metric based on labeled checkpoints for essential puzzles and game milestones. A visual comparison in Appendix˜E demonstrates the shortcomings of the original scores and shows how our metric provides a more representative signal of completion. The formal implementation of this metric is detailed in Section˜3.2.
@@
[Uncaptioned image] Kimi K2 10.5 8.3 19.7 9.0
[Uncaptioned image] GPT-4.1-mini 10.6 11.7 15.9 12.2
Table 1: LLMs performance on TextQuests. For complete results and more models, see Table˜3.
@@ 3.1 Evaluation Setting
The evaluation proceeds in a sequence of discrete turns. At each turn, the agent receives the latest observation from the environment, which is appended to the complete history of all previous observations, reasonings, and actions from the current game session. This full, multi-turn history is then provided as input to the model. The model’s task is to generate a brief reasoning for its strategy, followed by a single, executable command. Further details on the interaction protocol and system prompt are available in Appendix˜C.
For each model, we conduct two distinct evaluation runs: one with access to the game’s official clues (With Clues) and one without (No Clues). Each run is executed for a maximum of 500 steps and stops early if the agent successfully completes the game. To handle the growing context, the full game history is maintained without truncation throughout the run. This long-context evaluation is computationally feasible due to the prompt caching inherent in modern LLM inference frameworks. A detailed token analysis is provided in Table˜4.
Figure 2: Game progress for various models versus an optimal human walkthrough. Capable models sustain progress longer, suggesting better long-horizon reasoning.
@@ 4 Results and Analysis
Figure 3: Comparing mini and standard models from different closed-source providers
@@ 4.1 Quantitative Results
We evaluate a range of closed-source and open-weight models on TextQuests in two modes: with in-game hints (With Clues) and without (No Clues). As shown in Table˜1, even state-of-the-art LLMs make minimal progress in solving the games without assistance. In the With Clues setting, while access to the full hints allows all models to make more substantial progress, most still fail to complete the majority of the games. For instance, Sonnet 4 and Grok-3 each solved two games (Witness and Moonmist). Gemini 2.5 Pro and o3 solved these two and an additional game, Plunderedhearts. Opus 4 also solved Seastalker, bringing its total to four completed games. Furthermore, the performance differences between model sizes are large (Figure˜3), highlighting the importance of model scale for agentic tasks. This difficulty highlights that TextQuests is a challenging benchmark for measuring the long-horizon reasoning of LLM-based agents in exploratory environments.
@@ 4.2 Qualitative Analysis
To better understand why even capable models struggle with TextQuests, we analyze their trajectories to identify common failure modes. Figure˜4 illustrates common examples.
@@
Figure 5: A comparison of output and reasoning token efficiency across state-of-the-art LLMs on TextQuests. Since many exploratory steps are intermediate and don’t require a full reasoning budget, an ideal LLM agent should be efficient and dynamic with its reasoning effort while still maintaining consistent performance.
@@ A.1.2 Autosave
Appendix B Full Results
@@
Appendix C System Prompt and Environment Interaction
@@
As shown in Figure˜7, the evaluation proceeds in discrete history turns. For each turn, we parse the brief reasoning and corresponding action from the LLM’s output, adding them to the context history for the subsequent step. To ensure the agent’s decisions are based on its explicit plan, we discard any other ’thinking’ output and do not include it in the context history.
@@
Appendix D Token Analysis
We report the total input/output tokens cost to evaluate TextQuests in Table˜4.
@@ Appendix D Token Analysis
Appendix E Comparing Game Progress and Game Score
@@
To visually illustrate this discrepancy, Figure˜8 presents a direct comparison between the traditional Game Score and our checkpoint-based Game Progress metric. The figure highlights how our metric provides a more consistent signal of an agent’s approach to completion and shows clear cases where the game’s score is decoupled from this primary objective.

### v2

@@ 2 TextQuests
Previous work in text-based game evaluation has often relied on the games’ built-in scoring systems as the primary metric (Hausknecht et al., 2020; Yao et al., 2020). However, these point systems are a weak proxy for actual advancement, as they were often designed to reward exploration or enhance replayability rather than to track progress on the main storyline (for example, in The Witness, as many as 30 different endings are possible). To address these limitations, we introduce a new Game Progress metric based on labeled checkpoints for essential puzzles and game milestones. A visual comparison in Appendix˜F demonstrates the shortcomings of the original scores and shows how our metric provides a more representative signal of completion. The formal implementation of this metric is detailed in Section˜3.2.
@@
[Uncaptioned image] GPT-5 37.8 17.2 71.2 21.8
[Uncaptioned image] Claude Opus 4.1 33.9 19.1 68.0 22.1
@@
[Uncaptioned image] GPT-OSS 120B 12.0 21.2 18.1 13.8
Table 1: LLMs performance on TextQuests. All reasoning models are evaluated with high-reasoning budget. For complete results and more models, see Table˜4.
@@ 3.1 Evaluation Setting
The evaluation proceeds in a sequence of discrete turns. At each turn, the agent receives the latest observation from the environment, which is appended to the complete history of all previous observations, reasonings, and actions from the current game session. This full, multi-turn history is then provided as input to the model. The model’s task is to generate a brief reasoning for its strategy, followed by a single, executable command. Further details on the interaction protocol and system prompt are available in Appendix˜D.
For each model, we conduct two distinct evaluation runs: one with access to the game’s official clues (With Clues) and one without (No Clues). Each run is executed for a maximum of 500 steps, as we observed that longer runs yield only minimal additional game progress (see Appendix˜B). The run stops early if the agent successfully completes the game. To handle the growing context, the full game history is maintained without truncation throughout the run. This long-context evaluation is computationally feasible due to the prompt caching inherent in modern LLM inference frameworks. A detailed token analysis is provided in Table˜5.
Figure 2: Game progress for various model scales versus an optimal human walkthrough. Capable models sustain progress longer, suggesting better long-horizon reasoning.
@@ 4 Results and Analysis
Figure 3: Comparing mini and standard models from different closed-source providers, highlighting the importance of model scale for exploratory tasks.
@@ 4.1 Quantitative Results
We evaluate a range of closed-source and open-weight models on TextQuests in two modes: with in-game hints (With Clues) and without (No Clues). As shown in Table˜1, even state-of-the-art LLMs make minimal progress in solving the games without assistance. In the With Clues setting, while access to the full hints allows all models to make more substantial progress, most still fail to complete the majority of the games. Furthermore, the performance differences between model sizes are large (Figure˜3), highlighting the importance of model scale for agentic tasks. This difficulty highlights that TextQuests is a challenging benchmark for measuring the long-horizon reasoning of LLM-based agents in exploratory environments.
@@ 4.2 Qualitative Analysis
To understand why capable models are challenged by TextQuests, we analyze their trajectories to identify common failure modes.
@@
Figure 5: A comparison of output and reasoning token efficiency across state-of-the-art LLMs on TextQuests. Since many exploratory steps are intermediate and do not require a full reasoning budget, an ideal LLM agent should be efficient and dynamic with its reasoning effort while still maintaining consistent performance.
@@ A.1.2 Autosave
Appendix B Beyond 500 Run Steps
We extended our evaluation to $800$ steps for several models and observed that the increase in game progress was minimal after the 500-step mark. This trend is illustrated in Figure˜7, with detailed metrics provided in Table˜3.
Figure 7: Game progress trajectories on TextQuests for selected frontier LLMs evaluated up to 800800 steps. The visualization shows game progress saturating after approximately 500500 steps. For detailed metrics, see Table˜3.
No Clues With Clues
500 steps 800 steps 500 steps 800 steps
Progress (%) $\uparrow$ Harm $\downarrow$ Progress (%) $\uparrow$ Harm $\downarrow$ Progress (%) $\uparrow$ Harm $\downarrow$ Progress (%) $\uparrow$ Harm $\downarrow$
[Uncaptioned image] GPT-5 37.8 17.2 41.2 26.2 71.2 21.8 75.1 34.7
[Uncaptioned image] o3 30.9 18.7 35.7 25.4 60.4 17.2 63.2 26.1
[Uncaptioned image] Gemini 2.5 Pro 23.2 15.9 25.4 39.8 60.6 25.6 63.8 36.4
[Uncaptioned image] Claude Sonnet 4 24.7 16.0 27.4 26.9 57.2 18.4 58.6 27.6
[Uncaptioned image] GPT-5-mini 15.9 12.0 21.1 17.6 42.1 15.7 45.1 21.3
Table 3: Game progress and harm for several LLMs at 500500 and 800800 run steps.
Appendix C Full Results
@@
[Uncaptioned image] GPT-5 37.8 0 17.2 71.2 5 21.8
[Uncaptioned image] Claude Opus 4.1 33.9 0 19.1 68.0 4 22.1
@@
Appendix D System Prompt and Environment Interaction
@@
As shown in Figure˜8, the evaluation proceeds in discrete history turns. For each turn, we parse the brief reasoning and corresponding action from the LLM’s output, adding them to the context history for the subsequent step. To ensure the agent’s decisions are based on its explicit plan, we discard any other ’thinking’ output and do not include it in the context history.
@@
Appendix E Token Analysis
We report the total input/output tokens cost to evaluate TextQuests in Table˜5.
@@ Appendix E Token Analysis
Appendix F Comparing Game Progress and Game Score
@@
To visually illustrate this discrepancy, Figure˜9 presents a direct comparison between the traditional Game Score and our checkpoint-based Game Progress metric. The figure highlights how our metric provides a more consistent signal of an agent’s approach to completion and shows clear cases where the game’s score is decoupled from this primary objective.